Profitability ratios:
l. Profit margin %
m. Return on assets %
n. Return on equity %

SMOLIRA GOLF CORP.

2018 Income Statement

  Sales $ 336,329
  Cost of goods sold 231,000
  Depreciation 21,600
  Earnings before interest and taxes $ 83,729
  Interest paid 14,400
  Taxable income $ 69,329
  Taxes (21%) 14,559
  Net income $ 54,770
      Dividends $ 21,000
      Retained earnings 33,770


Some recent financial statements for Smolira Golf Corp. follow.

  

SMOLIRA GOLF CORP.
2017 and 2018 Balance Sheets
Assets Liabilities and Owners’ Equity
2017 2018 2017 2018
  Current assets   Current liabilities
      Cash $ 24,056 $ 24,200       Accounts payable $ 23,284 $ 27,200
      Accounts receivable 12,548 15,300       Notes payable 12,000 10,900
      Inventory 25,592 27,200       Other 11,671 15,900
        Total $ 62,196 $ 66,700         Total $ 46,955 $ 54,000
  Long-term debt $ 90,000 $ 93,294
  Owners’ equity
      Common stock and paid-in surplus $ 42,000 $ 42,000
      Accumulated retained earnings 208,936 242,706
  Fixed assets
  Net plant and equipment $ 325,695 $ 365,300   Total $ 250,936 $ 284,706
  Total assets $ 387,891 $ 432,000   Total liabilities and owners’ equity $ 387,891 $ 432,000

Find the following financial ratios for Smolira Golf Corp. (use year-end figures rather than average values where appropriate): (Enter your profitability ratio answers as a percent rounded to 2 decimal places, e.g., 32.16. Round the remaining answers to 2 decimal places, e.g., 32.16.)

  

j. Times interest earned times
k. Cash coverage ratio times
Long-term solvency ratios: 2017 2018
g. Total debt ratio times times
h. Debt-equity ratio times times
i. Equity multiplier times times
Short-term solvency ratios: 2017 2018
a. Current ratio times times
b. Quick ratio times times
c. Cash ratio times times
Asset utilization ratios:
d. Total asset turnover times
e. Inventory turnover times
f. Receivables turnover

times
Answer #1

Answer to Part a.

Current Ratio = Current Assets / Current Liabilities

2017:
Current Ratio = 62,196 / 46,955
Current Ratio = 1.32 times

2018:
Current Ratio = 66,700 / 54,000
Current Ratio = 1.24 times

Answer to Part b.

Quick Ratio = (Current Assets – Inventory) / Current Liabilities

2017:
Quick Ratio = (62,196 – 25,592) / 46,955
Quick Ratio = 36,604 / 46,955
Quick Ratio = 0.78 times

2018:
Quick Ratio = (66,700 – 27,200) / 54,000
Quick Ratio = 39,500 / 54,000
Quick Ratio = 0.73 times

Answer to Part c.

Cash Ratio = Cash and Cash Equivalents / Current Liabilities

2017:
Cash = 24,056 / 46,955
Cash Ratio = 0.51 times

2018:
Cash = 24,200 / 54,000
Cash Ratio = 0.45 times

Answer to Part d.

Total Asset Turnover = Sales / Average Total Assets
Average Total Assets = (387,891 + 432,000)/ 2
Average Total Assets = $409,945.50

Total Asset Turnover = 336,329 / 409,945.50
Total Asset Turnover = 0.82 times

Answer to Part e.

Inventory Turnover = Cost of Goods Sold / Average Inventory
Average Inventory = (25,592 + 27,200) / 2
Average Inventory = $26,396

Inventory Turnover = 231,000 / 26,396
Inventory Turnover = 8.75 times